In studying his campaign plans, Mr. Singleton wishes to estimate the difference between men's and women's views regarding his appeal as a candidate. He asks his campaign manager to take two random independent samples and find the 80% confidence interval for the difference. A random sample of 664 male voters and 711 female voters was taken. 245 men and 333 women favored Mr. Singleton as a candidate. Find this confidence interval.

Step 1 of 3:

Find the point estimate that should be used in constructing the confidence interval. Round your answer to three decimal places.

Step 2 of 3:

Find the margin of error. Round your answer to six decimal places.

Construct the 80% confidence interval. Round your answers to three decimal places.
Answer #1

Solution

P(Men voters favored Mr. Singleton as a candidate), Pm= 245/664 = 0.369

P(Women voters favored Mr. Singleton as a candidate), Pw = 333/711 = 0.468

point estimate = pm-pw = 0.369-0.468 =- 0.99

Margin of error : +-Z0.1 ((pm(1-pm)/nm) + (pw(1-pw)/nw))0.5

MOE : +- 1.28 ((0.369)(0.631)/664 + (0.468)(0.532)/711)0.5 = 0.033886

80% CI = ( - 0.99+- 0.033886) = (-1.024 , -0.956)